The following information applies to questions 19 - 24. Suppose we have the following information for the simple (fixed r, fi20. What is the at is the governments budget situation? (Hint: Think about what G laget situation? (Hint: Think about what
Answer #1

19).

Here the autonomous consumption is independent of disposable income. Here “400” is independent of disposable income, => autonomous consumption “400”. So, the correct answer is “A”.

20).

The government budget is given by, => B = T-G = 200 – 160 = 40 > 0, => the government budget is in surplus, => the correct answer is “B”.

21).

The equilibrium condition is given by.

=> Y = C+I+G, => Y = 400 + 0.9*(Y-200) + 300 + 160, => Y = 860 + 0.9*(Y-200).

=> 0.1*Y = 860 - 180, => Y = 680/0.1 = 6,800, => Y = $6,800 > 0.

The correct option is “B”.

22).

The consumption function is “C = 400 + 0.9*Yd”.

=> dC = 0.9*dYd, => dC = 0.9*400 = 360, => dC = 360 > 0. So, as the disposable income increases by “$400” the consumption increases by “$360”. The correct answer is “E”.